OpenStudy (anonymous):

Medal!!? AND FAN! please help someone! 22lbs. 14oz. 15lbs. 9oz. its additions. Please show work!

OpenStudy (anonymous):

@Jack1

OpenStudy (jack1):

um... what are we trying to do? are we trying to add the 2 weights together?

OpenStudy (anonymous):

yep:) sorry its not written better lol

OpenStudy (jack1):

all good, so there are 16 ounces (oz) in a pound (lb) so easiest thing to do is convert everything into ounces first, then add together, then convert back at the end to pounds... with me so far?

OpenStudy (anonymous):

yep got ya so far

OpenStudy (jack1):

sweet, 22 pound and 14 ounces is (22 x 16) + 14 = (352) + 14 = 366 ounces total

OpenStudy (jack1):

and 15 pounds 9 ounces = (15 x 16) + 9 = (240) + 9 = 249 ounces

OpenStudy (jack1):

so 366 + 249 = ...?@katlin95

OpenStudy (anonymous):

615!

OpenStudy (anonymous):

thanks you are the best!

OpenStudy (jack1):

perfect! now 615 / 16 = 38.44 pounds so 38 pounds and (0.44*16) = (7 ounces) so final answer would be 38 lb and 7 oz.

OpenStudy (jack1):

yw, slaters ;)

OpenStudy (triciaal):